OpenStudy (anonymous):

HELLLLLLLLLLLLLLLLLLP. The legs of a right triangle are 9cm and 11 cm long. find the length of the hypotenuse, approximated to three decimal places. a.14.213 cm b.12.876 cm c.8.998 cm d.21.902 cm

OpenStudy (anonymous):

use pythagoreans theorem knowing a=11,b=9 and a^2+b^2=c^2

OpenStudy (anonymous):

use pythagorean's theorem: A^2 +B^2 = C^2 where A and B are the lengths of the two sides and C is the length of the hypotenuse 9^2 + 11^2 = 202 sqrt(202) = 14.213

OpenStudy (anonymous):

a